The perimeter of a rectangle is 136 meters. The legnth is 2 meters longer than the width. Find the legnth and the width.
Formula being: P= 2L + 2W

Answer by Maths68(1474)

Length is two meters longer than the width means
l=w+2.......(1)
Put the value of length in give formula (P= 2L + 2W)
P=2l+2w
P=2(w+2)+2w
We know that P=136
136=2(w+2)+2w
136=2w+4+2w
or we can write
4w+4=136
4w=136-4
4w=132
w=132/4
w=33
Put the value of w in (1)
l=w+2.......(1)
l=33+2
l=35
Length is 35 and Width is 33 meters